Tom's father is 3times old as tom. 4years ago, he was 4 times older. how old is tom?
Answer by mananth(16946)

Let tom;s age be x
Tom's father will be 3x
4 years ago
tom was x-4 years
tom's father was 3x-4 years
four years ago 4times toms age = father's age
4(x-4)= 3x-4
4x-16=3x-4
4x-3x=16-4
x=12 which is tom's age.
Father's age will be 3*12
=36 years